OpenStudy (7thgrader):

What is the solution to the equation -2(1 + 5a) = -6(a+3) ? a =

OpenStudy (7thgrader):

@Kainui @Ashchu117 @sammixboo @LittleRedRiding @mathslover

OpenStudy (7thgrader):

@DuckDynastyfan923

OpenStudy (anonymous):

I would start by distributing to get rid of parenthesis, try that

OpenStudy (anonymous):

^^ exactly. can you do -2 * 1 and -2 * 5a?

OpenStudy (7thgrader):

so do i have to find out what number makes them the same number?

OpenStudy (anonymous):

ummmm... what?

OpenStudy (7thgrader):

-2 x 6a = -6(a + 3)

OpenStudy (anonymous):

you have to find out what a is equal to

OpenStudy (7thgrader):

-2 x 6 x a = -6 x (a + 3)

OpenStudy (anonymous):

I think what you mean is, you have to find a value for a so that both sides of the equation are equal, if thats what you mean, then yes thats true

OpenStudy (anonymous):

no, that is incorrect. @7thGrader your equation, after you distribute, would be -2 + -10a= -6(a+3)

OpenStudy (anonymous):

then you will distribute -6 to a and 3. -6 * a is -6a. -6 * 3 is what?

OpenStudy (7thgrader):

i know i need to learn but can someone get the answer for me? my mom said if i get this done i can stay home today

OpenStudy (7thgrader):

i will let you explain but will you give me what A equals first?

OpenStudy (7thgrader):

not to rush it's just i wanna stay home

OpenStudy (anonymous):

err, I'm going to have to walk you through it, just giving you the answer is against the CoC. can you do -6 * 3?

OpenStudy (7thgrader):

ye s

OpenStudy (7thgrader):

- 18

OpenStudy (anonymous):

good. now you have the equation -2 -10a = -6a - 18. next we will subtract -6a from both sides to get the variable on one side of the equation. after you do that, you have the equation -2 -6a -10a = -18.

OpenStudy (anonymous):

then you add 2 to both sides to cancel out the -2. what is -18 + 2?

OpenStudy (7thgrader):

-16

OpenStudy (anonymous):

good. now you have -6a -10a = -16. now we combine like terms- so what is -6 -10?

OpenStudy (7thgrader):

4?

OpenStudy (anonymous):

no, -16.

OpenStudy (anonymous):

now you have -16a = -16. divide both sides by -16 to get your answer. what is -16/-16?

OpenStudy (7thgrader):

you didn't put a + sign

OpenStudy (7thgrader):

1

OpenStudy (anonymous):

thats because everything is negative...

OpenStudy (anonymous):

correct! so you have your answer- a = 1. :)

OpenStudy (7thgrader):

thanks!

OpenStudy (anonymous):

no problem. :)

OpenStudy (anonymous):

(dont forget to close the question. )

OpenStudy (7thgrader):

can you help with one fast one i dont understand?

OpenStudy (7thgrader):

just did

OpenStudy (anonymous):

sorry, i cant right now, i gtg.

OpenStudy (anonymous):

i will maybe later tho

OpenStudy (ashchu117):

Why was I tagged here?